OpenStudy (anonymous):

Confirm that the Integral Test can be applied to the series. Then use the Integral Test to determine the convergence or divergence of the series. ln(2)/2 + ln(3)/3+ ln(4)/4+ ln(5)/5+ ln(6)/6+ . . . the problem is when using the intergal test its increasing from ln(2)/2 to ln(3)/3 and then decreses from onward....im not sure why its not working?

OpenStudy (anonymous):

@wio, this seems outta my league. Can you help this furry fellow out? Fur-rious I get it!

OpenStudy (anonymous):

additional info, from 1 to infinity and the formula can also be seen as ln(x+1)/(x+1) i know on the intergal test its positive, continues, but its not decreasing from the first term to the second term...thats the only problem i have with it.... its calc 2 any help :/

OpenStudy (anonymous):

The integral test would be showing that \[ \int_0^\infty \frac{\ln(x)}{x}dx \]converges, no? Are you having trouble with this integral?

OpenStudy (anonymous):

Or rather: \[\large \int_2^\infty \frac{\ln(x)}{x}dx \]Since we start at \(2\)

OpenStudy (anonymous):

Hmmm, right it needs to be monotonic.... My suggestion is to just just chop of the first term... If a series converges, adding a finite term to it doesn't make it stop converging.

OpenStudy (anonymous):

\[\large \sum_{n=2}^\infty \frac{\ln(n)}{n} =\frac{\ln(2)}{2} +\sum_{n=3}^\infty \frac{\ln(n)}{n} \]Since \(\ln(2)/2\) is finite, then we know either both series converge or both diverge.

OpenStudy (anonymous):

yup i got it on my own :) thanks for the answers though, i integrated it to confirm if it diverges which i got \[(\ln (x+1)^2)/2 \] and plugged in infinity and 1 for x and got it went to infinity, therefore \[\sum_{n=1}^{\infty}\]ln(x+1)/(x+1) diverges.